### v2

# SADMoE: Exploiting Activation Sparsity with Dynamic-$k$ Gating

Abstract
Transformer models, despite their impressive performance, often face practical limitations due to their high computational requirements. At the same time, such models exhibit significant activation sparsity, which can be leveraged to reduce the inference cost by transforming parts of the network into Mixture-of-Experts (MoE) layers. However, despite the crucial role of activation sparsity, its impact on this process remains unexplored. In this paper, we enhance the efficiency of MoE conversion through activation sparsity enforcement. Moreover, motivated by the high variance in the number of activated neurons, we propose a more effective dynamic-$k$ expert selection rule that adjusts the number of executed experts on a per-token basis. Finally, we extend this approach to multi-head attention projections, which results in even further savings. The proposed method, Sparsified Activation Dynamic-$k$ Mixture-of-Experts (SADMoE), outperforms existing approaches on common NLP and vision tasks, allowing us to save up to 60% of inference cost without significantly affecting model performance.
SADMoE: Exploiting Activation Sparsity with Dynamic-$k$ Gating
[…]
1 Introduction
footnotetext: ${}^{\ast}$ Equal contribution. Listing order is random. For a more detailed list of contributions please refer to Appendix Gfootnotetext: ${}^{\dagger}$ Correspondence to: bartwojc@gmail.com
Figure 1: Comparison of top-$k$ expert selection in standard MoE and dynamic-$k$ selection in SADMoE. With top-$k$, the same amount of compute is spent on each sample. dynamic-$k$ saves compute on easy inputs and can assign more experts to the hard ones to increase the performance.
Transformer has become a predominant model architecture in various domains of deep learning such as machine translation (Vaswani et al., 2017), language modeling (Devlin et al., 2018; Radford et al., 2019), and computer vision (Dosovitskiy et al., 2020; Khan et al., 2022). The widespread effectiveness of Transformer models in various applications is closely related to their ability to scale efficiently with the number of model parameters (Kaplan et al., 2020), prompting researchers to train progressively larger and larger models (Touvron et al., 2023; Jiang et al., 2023). However, the considerable computational demands of these models often restrict their deployment in practical settings with limited resources.
At the same time, Transformer models exhibit considerable activation sparsity in their intermediate representations Li et al. (2022), which suggests that most of their computations are redundant. Conditional computation methods reduce these unnecessary costs by using only a subset of the model parameters for any given input (Han et al., 2021). In particular, Mixture-of-Experts (MoE) layer (Shazeer et al., 2016), which consists of multiple experts that are sparsely executed for any single input, proved itself as an effective way to decouple the number of parameters of the model from its computational cost (Clark et al., 2022).
In this work, we aim to leverage MoE layers to decrease the inference cost in pre-trained dense models. As shown by Zhang et al. (2022), this can be done by converting dense layers into MoE layers. We demonstrate the relationship between the activation sparsity of the starting model and the efficiency of the final MoE model. We show that computational savings are directly related to sparsity levels, and propose to enforce higher activation sparsity levels before conversion, which leads to substantially improved cost-to-performance trade-off.
We further show that Transformer models exhibit significant variance in the number of activated neurons, which indicates that using standard top-$k$ expert selection is suboptimal. Therefore, we propose a dynamic-$k$ selection scheme (see Figure 1) that adjusts the number of activated experts on a per-token basis. To enable dynamic-$k$ selection, we propose a novel router training scheme that allows us to approximate the relative contribution of each expert, and select only a subset of the most relevant experts for inference. We find that this approach enables the model to efficiently allocate compute between easy and hard inputs, resulting in significant computational gains.
We extend this method to any standalone linear layer such as projections in Multi-Head Attention (MHA) layers, which often account for over 30%  (Song et al., 2022) of overall computations in the model. We propose a replacement procedure, in which every dense layer is substituted by a two-layer MLP module trained to imitate the output of the original layer. This conversion allows us to take advantage of activation sparsity in the replacement modules and further increase the computational gains.
Figure 2: Overview of key components of SADMoE: a) We propose to enhance the activation sparsity in the base model. b) We convert FFN layers in the model to MoE layers with routers that predict the expert contribution. c) We introduce dynamic-$k$ routing that selects the experts based on their relative contribution.
Our experiments on various NLP and vision benchmarks demonstrate that SADMoE achieves significant improvement in terms of cost-to-performance trade-off. Our contributions can be summarized as follows:
  • We demonstrate the impact of increased activation sparsity on the performance of dense-to-sparse-MoE conversion methods.
  • We propose a novel routing objective along with the dynamic-$k$ expert selection scheme that enables the model to select the appropriate number of experts for a given sample.
  • We extend the approach of dense-to-sparse-MoE to attention projections layers, which results in additional computational gains.
2 Related Work
Mixture-of-Experts.
Sparse Mixture-of-Experts layer (MoE) consists of a predefined number of experts along with a gating network that routes each input to a subset of experts. MoE layers were introduced as an efficient way to further increase the capacity of deep neural networks applied for NLP tasks, initially in LSTM models (Shazeer et al., 2016), and later in Transformers (Lepikhin et al., 2020). Since then, they have also been applied to computer vision (Riquelme et al., 2021; Daxberger et al., 2023). MoE layers have gained significant popularity primarily due to their excellent scaling properties (Du et al., 2022; Clark et al., 2022). Nonetheless, training such models is challenging, primarily because gating decisions must be discrete to ensure sparse expert selection. Various methods of training were proposed, some of which include reinforcement learning (Bengio et al., 2013), weighting the expert output by the probability to allow computation of the gradient of the router (Shazeer et al., 2016), or using the Sinkhorn algorithm (Clark et al., 2022). Some of those approaches also suffer from the possibility of load imbalance and therefore require auxiliary losses or alternative expert selection methods (Fedus et al., 2022; Zhou et al., 2022). Interestingly, in many cases fixed routing functions perform similarly to trainable routers (Roller et al., 2021), which suggests that current solutions are largely suboptimal. MoE models can also be derived from pre-trained dense models by splitting the model weights into experts and independently training the routers for each layer (Zhang et al., 2022; Zuo et al., 2022), which avoids most of the problems present in end-to-end training.
Activation sparsity in Transformers.
Li et al. (2022) show that ReLU-based Transformer models produce sparse activations in their intermediate representations, an effect that is prevalent across architectures, layers, and modalities. They propose a simple rule for keeping only top-$k$ activations in each MLP layer, which results in a model with comparable performance. Similarly, Mirzadeh et al. (2023) demonstrate that ReLU activation function in LLMs encourages ensuing activation sparsity that can be leveraged to skip redundant computations. Tuli and Jha (2023) take a step further and design a dedicated Transformer architecture accelerator that also exploits activation sparsity, while Liu et al. (2023) proposes to predict activation sparsity structure in LLMs and reduce the model latency by skipping redundant computations. Jaszczur et al. (2021) demonstrate that it is possible to train Transformer models from scratch with a fixed level of activation sparsity and obtain similar performance. Finally, a related line of works focuses on sparsity in the attention distributions instead of intermediate representations (Correia et al., 2019). None of the above-mentioned methods explore induced activation sparsity as a way to increase computational gains, nor do they address variance of the number of sparse activations on a per-token basis.
3 Method
SADMoE reduces the computational cost of the model by splitting every MLP module into a MoE layer. In this section, we describe all of its components in detail. The conversion procedure starts with activation sparsity enforcement, which we achieve by fine-tuning the pre-trained model with an auxiliary loss (Section 3.1). The weights in each block are then split into experts (Section 3.2), and the routers are trained to predict the norm of the output of each expert (Section 3.3). During inference, each MoE layer dynamically selects the number of experts to execute for any given sample following a simple dynamic-$k$ rule (Section 3.4). Finally, the whole process can be optionally preceded by MHA projections layer replacement, which allows us to apply the same conversion pipeline (Section 3.5) on the replacement modules. A high-level overview of the entire process is presented in Figure 2.
3.1 Enforcing activation sparsity
The scheme of reducing inference cost by dividing the model into independently activated modules relies on the well-known phenomenon of activation sparsity exhibited by most deep neural networks (Glorot et al., 2011). Taking inspiration from this observation, we expect that enforcing higher levels of activation sparsity may allow for the execution of an even smaller number of experts, resulting in overall computational savings. To this end, we induce activation sparsity by fine-tuning the model with an additional loss term that enforces higher levels of activation sparsity (Georgiadis, 2019). We apply the square Hoyer regularization (Kurtz et al., 2020; Hoyer, 2004) on the activations of the model:
$\mathcal{L}_{s}(x)=\frac{1}{L}\sum_{l=1}^{L}\frac{(\sum_{i}{|a^{l}_{i}|})^{2}}{\sum_{i}{{a^{l}_{i}}^{2}}},$ (1)
where $a^{l}$ is the activation vector from the middle layer of the $l$-th MLP for input $x$, and $L$ is the total number of MLPs in the model. Overall, the model is trained with the following cost function:
$\mathcal{L}(x)=\mathcal{L}_{\text{CE}}(\hat{y},y)+\alpha\mathcal{L}_{s}(x)$ (2)
where $\mathcal{L}_{\text{CE}}$ is cross-entropy loss, and $\alpha$ is the hyperparameter that controls the strength of sparsity enforcement.
3.2 Expert clustering
We split the two-layer MLP modules into experts using the parameter clustering method proposed by Zhang et al. (2022). Assuming the MLP layers are composed of weights $\bm{W}_{1}$, $\bm{W}_{2}$ and corresponding biases $\bm{b}_{1}$, $\bm{b}_{2}$, we treat the weights of each neuron from $\bm{W}_{1}$ as features and feed them into the balanced $k$-means algorithm (Malinen and Fränti, 2014) that groups neurons with similar weights together. Then, we use the resulting cluster indices to split the first linear layer $\bm{W}_{1}$, the first bias vector $\bm{b}_{1}$, and the second linear layer $\bm{W}_{2}$ into $n$ experts of the same size. The second bias $\bm{b}_{2}$ is not affected by this procedure.
3.3 Expert contribution routing
[…]
$\mathcal{L}_{r}(z)=\frac{1}{n}\sum_{i}^{n}{(R(z)_{i}-\lVert E_{i}(z)\rVert)^{2}}$ (3)
where $E_{i}$ is the $i$-th expert. We use a small two-layer neural network as the router $R$ and apply an absolute value activation function to ensure non-negative output. This regression-based formulation is still compatible with the commonly used top-$k$ expert selection, but enables more precise attribution of the contribution of each expert, as we show later in the experimental section.
Note that Zhang et al. (2022) also train each routing network independently, but their method constructs artificial labels for each input, and then subsequently trains the router as a classifier. We discuss the differences in detail in Appendix A.
3.4 Dynamic-$k$ gating
Figure 3: Distribution of non-zero activations in the FFN layers in GPT-2-base trained with ReLU on OpenWebText, with and without the sparsity enforcement phase. Both models exhibit significant variance, and the mean-to-variance ratio increases in the sparsified model.
Commonly used MoE layers always execute top-$k$ experts for each token, where $k$ is a predefined hyperparameter. This means that, regardless of the difficulty of the input, the model spends the same amount of compute on each batch (Zhou et al., 2022) or token (Shazeer et al., 2016). While this may be appropriate if the model is trained with the same restriction, we demonstrate why it is suboptimal for a model that was converted from a dense model without retraining. In Figure 3, we plot the distribution of non-zero activations in the intermediate layer of each FFN for a ReLU version of GPT-2, both for the standard model and a model trained with sparsity enforcement as described in Section 3.1. Activations in both models exhibit substantial variance, which indicates that the optimal number of executed neurons varies depending on the input token. Moreover, the sparsification phase increases the variance and further motivates per-token dynamic-$k$ expert selection.
[…]
$G(z)_{i}=\begin{cases}1&\text{if}\ R(z)_{i}\geq\tau\cdot\max R(z)\\ 0&\text{if}\ R(z)_{i}<\tau\cdot\max R(z)\end{cases}$ (4)
Note that as $\tau$ increases, the number of executed experts decreases along with the overall computational cost. We emphasize that after model deployment $\tau$ can be adjusted without the need for retraining.
3.5 Conversion of standalone dense layers
Figure 4: A large portion of the computational cost of Transformers is spent on linear projections in Multi-Head Attention. We replace each projection with a small MLP module, which allows us to apply our MoE conversion scheme on them. Attention heads are not shown for clarity of presentation.
A significant amount of computing in deep neural networks is often spent on dense layers that are not followed by any activation function. Dense-to-sparse-MoE conversion methods cannot reduce the costs of such layers due to lack of activation sparsity. This determines an upper bound on the possible computational savings. To overcome it, we substitute dense layers with small MLPs that approximately have the same computational cost and number of parameters. Each MLP is trained to imitate the output of the original dense layer given the same input by minimization of the mean squared error.
In our case, for Transformer architectures, we substitute projection matrices along with their biases in every MHA layer, as depicted in Figure 4. This means that the final model has four MoE layers in the MHA layer and one MoE layer in the FFN layer for each Transformer block. Note that we do not modify the computation of the scaled dot-product attention itself and that this scheme can be applied to any standalone dense layer.
Figure 5: FLOPs vs loss for GPT-2 on OpenWebText for MoEfication and SADMoE. Our method outperforms the baseline at every possible computational budget.
4 Experiments
To validate the impact of our method, we evaluate its performance on language modeling, text classification, and image classification. We obtain the performance versus computational cost characteristic for each method by evaluating the methods with different inference hyperparameters (either $\tau$ described in Section 3.4 for SADMoE or number of experts $k$ for MoEfication). We report the computational cost of each method in FLOPs [^1] We utilize the fvcore library to count model FLOPs: https://github.com/facebookresearch/fvcore, as it is a device-independent metric that has been shown to correlate well with latency (Mirzadeh et al., 2023).
For MoEfication, we follow the procedure described by Zhang et al. (2022) by converting the activation functions of the pre-trained model to ReLU and then fine-tuning the model. In the case of SADMoE, we also replace activation functions with ReLU, except for Section 5.2, where we demonstrate that our method performs well also with GELU. For the sake of a fair comparison, the total training data budget is always kept the same between different methods. See Appendix F for a more detailed description of the setup. The source code for our experiments is available at: https://github.com/bartwojcik/SADMoE.
4.1 Language modeling
Figure 6: Test accuracy of BERT models on CARER dataset at different compute levels. SADMoE displays significant improvements over MoEfication.
We evaluate our method on language modeling using GPT-2-base Radford et al. (2019) architecture by comparing it to MoEfication. We initialize the models from a publicly available OpenAI checkpoint pre-trained on a closed-source WebText dataset, and use OpenWebText (Gokaslan and Cohen, 2019) in all of our experiments. We train our model with sparsity enforcement loss on OpenWebText for 1000 steps, as we find out that it is sufficient for the model to recover the original loss value. We report the results in this Section without the MHA projection replacement step, as we find that GPT is especially sensitive to changes in attention layers, which always result in visible loss degradation.
Test losses for both methods are presented for different levels of compute budgets in Figure 5. Our method outperforms the baseline at every computational budget, owing its performance to the increased activation sparsity of the model and its ability to dynamically allocate the needed compute. Notably, the loss of SADMoE plateaus for higher budget levels, while the baseline displays consistently worse results whenever we lower the computational budget.
4.2 Text classification
We evaluate our method with BERT-base (Devlin et al., 2018) on the CARER dataset Saravia et al. (2018) that contains text samples categorized into 6 different emotion categories. We compare the accuracy versus FLOPs trade-off for SADMoE and MoEfication. For broader comparison, we also include a state-of-the-art early-exit method Zero-time Waste (ZTW) Wójcik et al. (2023). However, we should note that early-exit methods are orthogonal to MoE methods. The results are shown in Figure 6.
The performance of MoEfication gradually deteriorates and completely collapses when the number of executed experts approaches zero. In comparison, SADMoE maintains the original performance for a wide range of computational budgets, and the performance drop starts at a significantly lower budget. While early-exiting performs well for the lowest budgets, it obtains worse results than SADMoE at medium budgets and also suffers from gradual performance decline.
4.3 Image classification
Figure 7: ImageNet-1k FLOPs vs testset accuracy results for SADMoE, MoEfication and ZTW applied to ViT-B. Our method clearly outperforms the baselines on image classification due to elevated activation sparsity levels and dynamic-$k$ expert selection.
Vision Transfomers (Dosovitskiy et al., 2020) are becoming one of the most popular architectures in computer vision. Since our method is applicable to any Transformer model, we evaluate our method on the popular ImageNet-1k (Russakovsky et al., 2015) dataset. We use a pre-trained ViT-B checkpoint from the torchvision library [^2] https://pytorch.org/vision/stable/models.html as the base model and compare SADMoE with MoEfication and ZTW in terms of computational cost versus accuracy trade-off. Our results, presented in Figure 7, demonstrate the significant gains from applying our method over MoEfication, which emphasizes the role of induced activation sparsity and dynamic-$k$ expert selection rule.
5 Analysis
In this section, we present the results of experiments that provide insights into the performance of our method. Unless stated otherwise, we follow the experimental setup for models and datasets from the previous section.
Figure 8: Performance of ablated variants of SADMoE and MoEfication with BERT-base on the CARER dataset. We label the components of our method as S: sparsity enforcement, RR: regression routing, DynK: dynamic-$k$, and A: extension to attention projections. Each of the components of SADMoE meaningfully improves performance.
5.1 Ablation study
Since our method consists of several steps, the positive impact of each one of them may not be evident. To show the significance of every component, we perform an ablation study by incrementally adding each component to the baseline method. We take a BERT-base model and evaluate the ablated variants on CARER dataset in the same setting as described in Section 4.2. The results of this experiment are presented in Figure 8. As expected, each ablated version of the method improves upon the previous one. The sparsity enforcement phase leads to enhanced performance compared to plain MoEfication, as sparser activations enable better reconstruction of original output with fewer experts. Alternative router training objective and dynamic-$k$ expert assignment further improve the results, but – as the method only operates on the FFN layer – the computational cost cannot go below the cost of the remaining part of the model. Extending SADMoE to MHA projection layers allows our method to further reduce the computational cost, and the resulting full method retains the accuracy of the original model at around twice fewer FLOPs than MoEfication.
5.2 Sparsification and reliance on activation function
Figure 9: Comparison of our method applied on GPT-2 models with GELU and ReLU activation functions, with and without the activation sparsity enforcement phase. The sparsification enables the standard GELU model to perform similarly to the ReLU-based one, while the non-sparsified GELU model is inferior to its ReLU counterpart.
Activation sparsity works focus their analysis on networks with ReLU, as activation functions commonly used in Transformer models (such as GELU or SiLU) do not output exact zeros for inputs below zero. Although activations with near-zero values might have a negligible impact on the final output, assessing their number is non-trivial. They also make it hard to apply activation sparsity enforcement schemes, such as $\ell^{1}$ penalty, that mostly assume the ReLU activation function. To get around this limitation, the works on activation sparsity usually change the activation function of the trained model to ReLU and continue pre-training the modified model until it reaches the performance of the original model  (Zhang et al., 2022; Mirzadeh et al., 2023).
Figure 10: Effect of expert division granularity on our method. In general, smaller experts display favorable performance as they allow for larger computational savings.
We hypothesize that the dependence on ReLU is not necessary and that the models with many near-zero activations can effectively function in a very similar way to standard activation-sparse ReLU-based models. To evaluate this hypothesis, we extend the sparsity enforcement scheme to the GELU activation function by penalizing the model for pre-activation values larger than a certain threshold. We first transform the pre-activation values:
$z^{\prime}=\max(0,z-d),$ (5)
where $z$ is the pre-activation value and $d$ is the displacement hyperparameter. Then, we apply the loss from Equation 1 on $z^{\prime}$. The purpose of this transformation is to penalize the model only for pre-activation values larger than $d$. We empirically find that $d=-10$ works well with GELU activation, as the output of GELU activation function below this value is near-zero. Following the methodology from Section 4.1, we train ReLU-based and GELU-based GPT-2 with and without sparsity enforcement loss and show the results of our experiment in Figure 9. SADMoE with a sparsified GELU-based model performs similarly to a sparsified ReLU-based model, while the performance of the non-sparsified GELU-based variant rapidly collapses for $\tau>0$. Within ReLU-based models, the sparsification still enhances the performance of SADMoE, but the improvements are less drastic, and the behavior of our method does not significantly change as in the case of GELU. This shows sufficient activation sparsity enforcement relieves the model from the dependence on ReLU.
Figure 11: Per-layer distribution of the number of executed experts in SADMoE trained on the CARER with different $\tau$ thresholds for a standard, non-sparsified model (top row) and a sparsified model (bottom row). The high variability of that number explains the computational gains from using dynamic-$k$.
5.3 Impact of expert granularity
A crucial hyperparameter in SADMoE is the selection of expert size. Smaller experts may allow for a more granular selection of executed neurons, thus likely resulting in a lower computational cost. However, decreasing the expert size increases the number of experts, which translates to a larger router, potentially negating any computational gains. To study the impact of this hyperparameter on our method, we evaluate SADMoE on GPT-2 with experts of different sizes, and show the results in Figure 10.
We observe that our method generally performs better with smaller experts. Those results differ from the results presented in  (Zhang et al., 2022), where the expert size is significantly higher. The positive correlation between granularity and performance can be explained by the increased levels of activation sparsity in our model, which requires significantly fewer activated neurons (experts) to correctly function. As expected, the performance decreases for the extreme choice of expert size equal to 1 due to significantly larger routing costs. We include additional results for expert granularity with GELU activation in Appendix D.
5.4 Expert selection patterns
To explore the variability in SADMoE expert usage patterns, we investigate the distribution of the number of executed experts for different values of the $\tau$ threshold, with and without the sparsification phase. In Figure 11, we show the histograms of the numbers of activated experts for each layer FFN of the BERT-base trained on the CARER dataset. As expected, the model with enforced activation sparsity requires a smaller number of experts for each given threshold threshold. Deeper model layers generally utilize more experts, which suggests that the features at those levels are more complex. Interestingly, even for high values of $\tau$, the sparsified model does not always use all available experts. Both base and sparsified models also exhibit significant variance in the number of activated neurons across different layers, which justifies the dynamic-$k$ selection and indicates that computational adaptability mechanisms are crucial for efficient inference in Transformer-based models.
6 Conclusion
We introduce Sparsified Activation Dynamic-$k$ Mixture-of-Experts (SADMoE), a novel approach that induces activation sparsity to improve the efficiency of Transformer-based models by converting their layers to Mixture-of-Experts (MoE). We demonstrate the interplay between the activation sparsity of dense models and the efficiency of converted MoE. Moreover, we introduce regression-based router training and dynamic-$k$ routing, which enable our method to efficiently utilize the induced sparsity. Finally, we show how dense-to-sparse-MoE conversion approaches can be extended to standalone dense layers. Our approach is compatible with existing Transformer architectures like GPT, BERT, and ViT, and significantly improves upon the existing MoE conversion schemes. Our findings contribute to the ongoing efforts to make Transformer models more efficient and accessible for a wider range of applications, especially in resource-constrained environments.
Limitations
While SADMoE displays promising results in reducing the computational cost of inference in Transformer models, several limitations should be acknowledged. Our proposed sparsity enforcement and router training phases require additional training time. This overhead, while small, should be considered when evaluating the overall efficiency gains of our approach. Moreover, we demonstrate improved performance over existing approaches on common NLP and vision tasks, but the scope of our experiments is restricted due to limited access to computational resources. Further research is needed to address the aforementioned limitations and to explore its applicability to extremely large models. We focus primarily on fundamental machine learning research and we do not see any specific risks or ethical issues associated with our method. Nevertheless, we recognize the potential for misuse of machine learning technology and advocate for responsible AI practices to mitigate such risks.
Acknowledgments
We gratefully acknowledge Poland’s high-performance Infrastructure PLGrid (HPC Centers: ACK Cyfronet AGH, PCSS, CI TASK, WCSS) for providing computer facilities and support within computational grants no. PLG/2023/016393 and PLG/2023/016321. Filip Szatkowski is supported by National Centre of Science (NCP, Poland) Grant No. 2022/45/B/ST6/02817. Bartosz Wójcik is supported by National Science Centre, Poland, Grant No. 2023/49/N/ST6/02513.
Appendix A Difference between router training in SADMoE and MoEfication
Our router training procedure might seem similar to the one proposed in MoEfication (Zhang et al., 2022), but the source code of the method provided by the authors [^3] MoEfication source code for router training is publicly available at: https://github.com/thunlp/MoEfication/blob/c50bb850307a36f8a0add6123f56ba309a156d13/moefication/utils.py#L188-L260 contains a different routing objective than the one reported in the paper. While the paper describes their router training objective as a prediction of the sum of ReLU activation values in each expert, the source code uses prediction labels created from the sum of the activations in the intermediate layer divided by the maximum value in the entire batch and minimize the binary cross-entropy loss. Assuming that $a_{k,j}$ is the activation vector in the hidden layer of expert $j$ for sample $k$, the label generation for their router can be expressed as:
$y_{k,j}=\frac{\sum_{i}a_{k,j,i}}{\max_{l,m}\sum_{i}a_{l,m,i}}$ (6)
In comparison to their approach, the router training procedure in SADMoE differs in multiple aspects:
  • Our router considers the output of each expert instead of looking at the activations in the intermediate layers.
[…]
  • Our router is trained with the mean-squared error instead of the binary cross-entropy loss. The output of our router is constrained to positive values, while the MoEfication router is constrained to outputs in $[0,1]$.
We find that the above differences are responsible for the improved performance of our router (see Figure 8).
Appendix B Comparision of FLOPs between standard FFN layer and dynamic-$k$ MoE
To compare the efficiency of our method with a standard MLP layer in Transformer, we estimate FLOPs in both modules. We assume the layer is composed of two linear transformations, with input and output size $d_{m}$ and hidden dimension $ed_{m}$, where $e$ is the expansion factor, which is usually equal to 4 in standard Transformer models. We skip the negligible cost of the biases and activation functions for simplicity.
One can estimate the cost of the MLP layer in FLOPs, $C_{\text{MLP}}$, as:
$C_{\text{FFN}}=d_{m}\cdot ed_{m}+ed_{m}\cdot d_{m}={d_{m}}^{2}\cdot 2e.$ (7)
For dynamic-$k$ expert selection with $n$ total experts and $k$ experts selected for a given input, the cost of the forward pass is composed of the cost of a forward pass through $k$ experts and the cost of the 2-layer router with input dimension $d_{m}$, hidden dimension $d_{h}$ and output dimension $n$. The cost of the single expert pass can be expressed as:
$C_{E}=(d_{m}\cdot\frac{ed_{m}}{n}+\frac{ed_{m}}{n}\cdot d_{m})={d_{m}}^{2}\cdot\frac{2e}{n},$ (8)
and the routing cost can be estimated as:
$C_{R}=d_{m}\cdot d_{h}+d_{h}\cdot n.$ (9)
Therefore, the full cost of dynamic-$k$ $C_{\text{dynk}}$ can be estimated as:
$C_{\text{dynk}}=k\cdot C_{E}+C_{R}={d_{m}}^{2}\cdot\frac{2ek}{n}+d_{h}(d_{m}+n),$ (10)
and the cost of our method compared to the cost of standard MLP can be expressed as:
$\frac{C_{\text{dynk}}}{C_{\text{MLP}}}$ $=\frac{{d_{m}}^{2}\cdot\frac{2ek}{n}+d_{h}(d_{m}+n)}{{d_{m}}^{2}\cdot 2e}$ (11)
$=\frac{k}{n}+\frac{d_{h}(1+\frac{n}{d_{m}})}{{d_{m}}\cdot 2e}.$ (12)
As long as the number of selected experts $k$ does not approach the total number of experts $n$ and the hidden dimension of the router does not approach the size of hidden dimension $d_{m}$, the ratio is significantly below one.
Assuming the worst case for second term ($n=ed_{m}$), we can estimate the cost ratio as:
$\frac{k}{n}+\frac{d_{h}}{d_{m}}\cdot\frac{1+e}{2e},$ (13)
which shows that dynamic-$k$ expert selection only exceeds the FLOPs cost of the standard network when the dynamic-$k$ rule selects almost all experts or the number of experts becomes very high. For an even more detailed analysis, we refer to Figure 14 where the the cost ratio between our method and standard MLP is shown, assuming different router sizes and $e=4$ as standard for most Transformer models. In practice, we use $d_{h}=128$, so in all our experiments $d_{m}=6d_{h}$.
Appendix C Base model activation sparsity
In this section, we investigate the impact of activation sparsity levels on the performance of our method. We train multiple models with different activation sparsity enforcement loss weights $\alpha$ and convert them to Mixture-of-Experts with our method and MoEfication. Note that, in contrast to other sections in this work, we apply MoEfication to the model fine-tuned with sparsity enforcement.
The results highlight the positive correlation between the activation sparsity and the performance of the converted MoE, as higher sparsity in the base model always translates to better performance for both SADMoE and MoEfication. Our method achieves better results for every base model in all cases, which proves that dynamic-$k$ selection makes better use of the induced sparsity.
Figure 12: FLOPs vs loss for GPT-2 on OpenWebText for SADMoE (left) and MoEfication (right) with different sparsity levels in the base model. Sparsity enforcement is beneficial for both methods, but our one outperforms MoEfication baseline as dynamic-$k$ can better utilize the sparsity.
Appendix D Additional results with expert size and GELU
Figure 13: Impact of expert division granularity on our method with the GELU activation function.
In addition to experiments in Section 5.3, we present the results of similar ablation carried on the sparsified GPT-2 model with GELU activation. The results, presented in Figure 13, follow the same pattern as before, which supports our claim that the sparsification enables the GELU-based models to function similarly to ReLU-based ones.
Appendix E Expert activation patterns for attention projection layers
Following the analysis for MoE-converted FFN layers in Section 5.4, we investigate the activation patterns in MHA projections modified with our method in Figures 15, 16, 17 and 18. The projection modules display lower levels of sparsity than FFNs, which is to be expected as our projection layers have lower intermediate dimensionality. Expert selection distribution patterns in $Q$ and $K$ show significant similarity, and the patterns in $V$ and output projections are also similar to a lesser degree. The variance in the number of selected experts in MHA projections is higher than in FFN layers, but it still exists and the distribution in some of the layers seems to be bimodal, which still justifies dynamic-$k$ selection rule.
Appendix F Training details
In this Section, we describe the technical details used in the SADMoE conversion procedure. For full reproducibility, we also share the source code that we used for conducting the experiments. All experiments were performed using the PyTorch library (Paszke et al., 2019) on the NVIDIA A100 and V100 GPUs.
F.1 Language modeling
We base our code and hyperparameters on the nanoGPT repository provided at https://github.com/karpathy/nanoGPT. In all pretraining experiments, we keep the fixed number of steps at 1000 and keep the effective batch size equal to the value in the repository through gradient accumulation. The alpha values for sparsity enforcement can be found at Figure 12. We train the routers for SADMoE and MoEfication for 2000 steps using one GPU and tuning the learning rates for a given expert size from the range between $0.002-0.005$. For router training, we use Adam optimizer and cosine warmup scheduler.
F.2 Text classification
All experiments start from the same pre-trained BERT-base checkpoint. For methods requiring ReLU activation function, we replaced GELU with ReLU and continue model pretraining on concatenated wikipedia (Foundation, ) and books (Zhu et al., 2015) corpora for $5000$ steps on $8$ GPUs using main setup from https://github.com/huggingface/transformers/blob/main/examples/pytorch/language-modeling/run_mlm.py, per device batch size $96$ and learning rate $5\cdot 10^{-4}$. For MHA projections replacement we use the same corpus and train replaced MLP modules on a single GPU with batch size $128$ and learning rate $0.001$ for $3000$ steps. We finetuned base dense models on CARER datasetw for $5$ epochs with $2\cdot 10^{-5}$ learning rate. For sparsity enforcement in SADMoE we use $\alpha$ linearly increasing from zero to $0.001$ over training. For both MoEfication and SADMoE we train routers with batch size 64 and initial learning rate $0.001$ for 5 epochs. In all experiments, we use Adam optimizer with linear learning rate decay. For MoEfication we use same expert size $32$, for SADMoE we use more granular expert size equals $6$. For ZTW we trained ICs for 5 epochs with batch size $32$ and learning rate $0.01$.
F.3 Image classification
All methods start with the same pre-trained ViT-B from the torchvision library and are trained on ImageNet-1k using the augmentation proposed by Touvron et al. (2022). We use mixup ($0.8$), cutmix, label smoothing ($0.1$), gradient clipping ($1.0$) and the Adam optimizer with cosine learning rate schedule without warm-up. For SADMoE, we replace the MHA projections and train the replacements for $3$ epochs with the initial learning rate $0.001$ and batch size $128$, and then finetune the model for $90$ epochs with sparsity enforcement weight $\alpha=0.2$, initial learning rate $2\cdot 10^{-5}$ and batch size $512$. We then convert the modules into MoE layers, and train the gating networks for $7$ epochs with the initial learning rate set to $0.001$ and batch size $128$. We train ZTW for $100$ epochs in total, allocating $5$ epochs for ensemble training, while keeping the rest of the original hyperparameters unchanged. For MoEfication, we first convert the pre-trained model to ReLU-based one and finetune for $90$ epochs with an initial learning rate of $0.0001$ and batch size $256$. We then split the weights and train the routers for $10$ epochs with the initial learning rate $0.001$ and batch size $256$.
Appendix G Contributions
Filip integrated GPT-2 into the codebase and ran GPT-2 experiments, performed the distribution of non-zero activations analysis, the expert granularity analysis, and the activation function analysis. He contributed to the design of dynamic-k gating, and played a primary role in designing the experiments, writing of the article and producing its visualizations.
Bartosz set the research direction of the project and proposed the alternative routing scheme, dynamic-k expert selection, the additional activation sparsity enforcement phase for ReLU and GELU, and the replacement of MHA projection layers. He also wrote the shared codebase for the experiments, carried out the ViT-B experiments, and also played a primary role in the writing and editing of the article.
Mikołaj made this paper possible by performing all of the experiments at the initial stages of the project and implementing MoEfication and numerous variants of our method. He carried out the BERT experiments, performed the expert selection pattern analysis and the ablation study, and contributed to the crafting of the paper.
Kamil helped with the writing and editing of the paper.
Figure 14: FLOPs ratio between dynamic-$k$ expert layer and standard two-layer MLP for different values of the total number of experts $n$ and number of selected experts $k$. We assume the hidden dimension of router $d_{h}$ is based on model dimension $d_{m}$, and set standard expansion factor $e=4$. For different sizes of router, dynamic-$k$ uses fewer FLOPs than standard MLP as long as the total number of experts is sufficiently large and the number of selected experts is not equal to the total number of experts. For the clarity of presentation, we plot discrete values of $k$ and $n$ as continuous.
Figure 15: Distribution of the number of executed experts in each layer for query projections.
Figure 16: Distribution of the number of executed experts in each layer for key projections.
Figure 17: Distribution of the number of executed experts in each layer for value projections.
Figure 18: Distribution of the number of executed experts in each layer for output projections.